OpenStudy (anonymous):

Please help medal!!! A bakery formula developed by Awesome Bread calls for 6 1/2oz. of baking soda for 200 loaves. Joyce wants to adapt the recipe to go into buisness for herself. but she only wants to make 30 loaves. How much baking soda will she need?

OpenStudy (kittiwitti1):

So, 6.5 oz baking soda per 200 loaves Put it in fraction form as a ratio\[\frac{6.5}{200}\]

OpenStudy (anonymous):

=0.0325

OpenStudy (kittiwitti1):

No, no simplifying. It's just a ratio. The problem is asking for how many oz. of baking soda we need for THIRTY loaves... Proportion:\[\frac{6.5}{200}=\frac{x}{30}\]Do you know how to solve a proportion?

OpenStudy (anonymous):

find the common denomenator right?

OpenStudy (kittiwitti1):

Well... I don't know about that method, but I would cross-multiply.

OpenStudy (anonymous):

or we can use 1 for x right?

OpenStudy (anonymous):

so cross multiplying i got 195

OpenStudy (anonymous):

195/30 =6.5 oz

OpenStudy (kittiwitti1):

\[200x=30(6.5)\text{, isolate the x}\]